# Singularities of base loci on abelian varieties

### v1

@@ Abstract.
We prove that the log canonical threshold of the base ideal of a complete linear system on an abelian variety is $\geq 1$, and equality holds if and only if the base locus has divisorial components.
@@
A well known theorem of Kollár ([K1, Theorem 17.13]) asserts that for a principally polarized abelian variety the pair $(A,\Theta)$ is log-canonical. This simple and fundamental result has been subsequently sharpened by Ein and Lazarsfeld in [EL], and there are still important open conjectures in this field, see e.g. [CGS], [MP, §H.29], [SY, §9] and references therein. In turn the theorems of Kollár and Ein-Lazarsfeld have been extended to other integral and/or $\mathbb{Q}$-divisors on abelian varieties in [H1],[H2],[DH],[P], [B], [JL].
In this note Kollár’s result is extended in a different direction: base loci of complete linear systems. For an effective line bundle $L$ on a complex abelian variety $A$ we denote $\mathfrak{b}_{L}$ the base ideal sheaf of the complete linear system $|L|$, namely the image of the twisted evaluation map $H^{0}(A,L)\otimes L^{-1}\rightarrow{\mathcal{O}}_{A}$. Our result is
@@
(When there are divisorial components the situation is completely understood, see the last part of the proof below). When there are no divisorial components in the base locus we don’t expect our result to be close to sharp, and a better understanding of the log canonical threshold (and more generally of the jumping numbers) of base ideals on abelian varieties seems to be an interesting and difficult problem.
@@
The result of this paper was obtained while I was working with Enrico Arbarello and Giulio Codogni on the reducedness of the base scheme of line bundles on abelian varieties, and on the other topics of [ACP]. I thank them for their generosity, and for valuable conversations and suggestions. More recently, I have greatly benefitted of a continuous exchange with Arbarello.
@@ 2. The proof
Let $L$ be an ample line bundle on an abelian variety $A$. Let $\mathcal{I}\subset{\mathcal{O}}_{A}$ be a $K(L)$-invariant ideal sheaf, i.e. $t_{x}^{*}\mathcal{I}=\mathcal{I}$ for all $x\in K(L)$. Then, for $y\in A$, either $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=H^{0}(A,t_{y}^{*}L)$ or $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=0$. Moreover if $\mathcal{I}$ is non-trivial then the Zariski open set of $y\in A$ such that $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=0$ is non-empty.
@@ 2.2. Proof of the Theorem
Conversely, if $L$ is ample and $\mathfrak{b}_{L}$ has some divisorial components it is known that $A\cong B\times C$, where $B$ is a principally polarized abelian variety, $C$ is an abelian variety, and $L={\mathcal{O}}_{B}(\Theta_{B})\boxtimes L_{C}$, where $L_{C}$ is an ample line bundle on $C$ without base divisors ([BL, Theorem 4.3.1]). It follows that $\mathfrak{b}_{L}={\mathcal{O}}_{B}(-\Theta_{B})\boxtimes\mathfrak{b}_{L_{C}}$. Hence, by the above, $\mathcal{J}(\mathfrak{b}_{L})=p_{B}^{*}{\mathcal{O}}_{B}(-\Theta_{B})$.

### v2

@@ Abstract.
We prove that the log canonical threshold of the base ideal of a complete linear system on a complex abelian variety is $\geq 1$, and equality holds if and only if the base locus has divisorial components. Consequently the same assertions hold for the ideal of the intersection of translates of theta divisors by the points of a finite subgroup.
@@
A well known theorem of Kollár ([K1, Theorem 17.13]) asserts that for a complex principally polarized abelian variety the pair $(A,\Theta)$ is log-canonical. This simple and fundamental result has been subsequently sharpened by Ein and Lazarsfeld in [EL], and there are still important open conjectures in this field, see e.g. [CGS], [MP, §H.29], [SY, §9] and references therein. In turn the theorems of Kollár and Ein-Lazarsfeld have been extended to other integral and/or $\mathbb{Q}$-divisors on abelian varieties in [H1],[H2],[DH],[P], [B], [JL].
In this note Kollár’s result is extended in a different direction: base loci of complete linear systems, and intersections of translates of theta divisors by the points belonging to a finite subgroup. We work over $\mathbb{C}$. For an effective line bundle $L$ on an abelian variety $A$ we denote $\mathfrak{b}_{L}$ the base ideal sheaf of the complete linear system $|L|$, namely the image of the twisted evaluation map $H^{0}(A,L)\otimes L^{-1}\rightarrow{\mathcal{O}}_{A}$. Our result is
@@
(When there are divisorial components the situation is completely understood, see the last part of the proof below). By a standard argument it follows that the conclusion of Theorem holds also for the (scheme-theoretic) intersection of translates of theta divisors of a principally polarized abelian variety $A$ by all points of a finite subgroup $H\subset A$.
Here and in the sequel, given a morphism $f:X\rightarrow Y$ and an ideal sheaf $\mathfrak{a}$ on $Y$ we will denote $f^{-1}\mathfrak{a}$ the inverse image ideal sheaf on $X$, i.e. the ideal sheaf generated by the pullbacks of local sections of $\mathfrak{a}$.
Corollary.
Let $A$ be a principally polarized abelian variety, $\Theta$ a theta divisor, and $H$ a finite subgroup of $A$. Let
$\mathfrak{a}_{\Theta,H}=\sum_{h\in H}t_{h}^{-1}{\mathcal{O}}_{A}(-\Theta)\subset{\mathcal{O}}_{A}.$
Then
$lct(\mathfrak{a}_{\Theta,H})\geq 1$
and equality holds if and only if the zero locus of $\mathfrak{a}_{\Theta,H}$ has divisorial components.
When there are no divisorial components we don’t expect our results to be close to sharp, and a better understanding of the log canonical threshold (and more generally of the jumping numbers) of base ideals on abelian varieties seems to be an interesting and difficult problem.
@@
The result of this paper was obtained while I was working with Enrico Arbarello and Giulio Codogni on the reducedness of the base scheme of line bundles on abelian varieties, and on the other topics of the erratum in [ACP]. I thank them for their generosity, and for valuable conversations and suggestions. More recently, I have greatly benefitted of a continuous exchange with Arbarello.
@@ 2. Proofs
Let $L$ be an ample line bundle on an abelian variety $A$. Let $\mathcal{I}\subset{\mathcal{O}}_{A}$ be a $K(L)$-invariant ideal sheaf, i.e. $t_{x}^{-1}\mathcal{I}=\mathcal{I}$ for all $x\in K(L)$. Then, for $y\in A$, either $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=H^{0}(A,t_{y}^{*}L)$ or $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=0$. Moreover if $\mathcal{I}$ is non-trivial then the Zariski open set of $y\in A$ such that $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=0$ is non-empty.
@@ 2.2. Proof of the Theorem
Conversely, if $L$ is ample and $\mathfrak{b}_{L}$ has some divisorial components it is known that $A\cong B\times C$, where $B$ is a principally polarized abelian variety, $C$ is an abelian variety, and $L={\mathcal{O}}_{B}(\Theta_{B})\boxtimes L_{C}$, where $L_{C}$ is an ample line bundle on $C$ without base divisors ([BL, Theorem 4.3.1]). It follows that $\mathfrak{b}_{L}={\mathcal{O}}_{B}(-\Theta_{B})\boxtimes\mathfrak{b}_{L_{C}}$. Hence, by the above, $\mathcal{J}(\mathfrak{b}_{L})=p_{B}^{*}{\mathcal{O}}_{B}(-\Theta_{B})$. This concludes the proof of the Theorem.
2.3. Proof of the Corollary
We prove more generally the following
Corollary 2.3.1.
Let $L$ be an effective line bundle on an abelian variety and let $H$ be a finite subgroup of $A$. Then
$lct(\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L})\geq 1$
and equality holds if and only if the zero scheme of $\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L}$ has some divisorial components.
Proof.
The argument is standard and similar to [ACP, Corollary 8]. Let $\varphi_{L}:A\rightarrow{\rm Pic}^{0}A$ be the homomorphism associated to $L$ and let $K:=\varphi_{L}(H)$. Let us consider the isogeny
$f:{\rm Pic}^{0}A\rightarrow({\rm Pic}^{0}A)/K$
and let
$\pi_{K}:B\rightarrow A$
be the dual isogeny. Then
$H^{0}(B,\pi_{K}^{*}L)\cong\bigoplus_{\alpha\in K}\pi_{K}^{*}H^{0}(A,L\otimes P_{\alpha}).$
Therefore
$\mathfrak{b}_{\pi_{K}^{*}L}=\pi_{K}^{-1}(\sum_{\alpha\in K}\mathfrak{b}_{L\otimes P_{\alpha}})=\pi_{K}^{-1}(\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L}).$
Corollary 2.3.1 follows from the Theorem because $\mathcal{J}(\mathfrak{b}_{\pi_{K}^{*}L})=\pi_{K}^{-1}\mathcal{J}(\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L})$ ([L, Example 9.5.44]). ∎